OpenStudy (anonymous):

i guess this isnt really a feedback, however it is a question about openstudy in general. what do the different colors your name is in mean (i know purple is mod,) and what are the purpose of the shapes that have your smart score in them? some are diamonds and others are circles. the ocd has gotten to me haha

sammixboo (sammixboo):

Well, the different colors in the username shows your level of SmartScore 0-19 = Red 20-49 = Orange 50-74 = Yellow 75-89 = Olive 90-100 = Light Green A regular user who hasn't subscribed to OpenStudy gets a circle around their SmartScore, but if you have subscribed then you get a diamond around your SmartScore! Do you have any other questions?

OpenStudy (anonymous):

ahhh, that makes a little bit more sense. what is subscribing/ how do you do it/ whats the purpose?

sammixboo (sammixboo):

Subscribing is donating money to OpenStudy to help improve the site :) If you go to your settings and then go to the Account Settings, you can see what is Subscription and it's purpose :) You can also subscribe from there

OpenStudy (anonymous):

ahhh. im all caught up for now. ty(:

sammixboo (sammixboo):

No problem!